OpenStudy (anonymous):

how to evaluate 8^-4

ganeshie8 (ganeshie8):

\(we\ have: \ \huge a^{-m} = \frac{1}{a^m}\)

ganeshie8 (ganeshie8):

\(\huge 8^{-4} = \frac{1}{8^4} = \frac{1}{4096}\)

OpenStudy (anonymous):

oh thanks ,

ganeshie8 (ganeshie8):

welcome.. :))

OpenStudy (unklerhaukus):

\[8^{-4}=\frac18\times\frac18\times\frac18\times\frac18\]